OpenStudy (anonymous):

Approximate using linearization and compute percentage error: 1/(10.03)^2

OpenStudy (anonymous):

let \[f(x)=\frac{ 1 }{ x^2 } , x1=10 , dx=x-x1=.03\] \[y=y1+f \prime(x1)*dx\]\[y1=\frac{ 1 }{ x1^2 }=\frac{ 1 }{ 10^2 }=.01\] \[f \prime(x1)=\frac{ -2 }{ x1^3 }=\frac{ -2 }{ 10^3 }=-.002\]\[y=.01-002*.03\]

OpenStudy (anonymous):

so \[\frac{ 1 }{ 10.03^2 }=.01-.0006=.0094\]

OpenStudy (anonymous):

Hey, you are the man. Thank you so much!

OpenStudy (anonymous):

you are so welcome :) , and I am sorry if I got any wrong in these calculations

OpenStudy (anonymous):

hold on, when you say x1, would that also mean f(a)? my professor seems to use different terms than everyone else

OpenStudy (anonymous):

No worries, the concept was my only concern, I can double check the calculations. Thanks!

OpenStudy (anonymous):

no no yuor prof is right x1,a, they are just numbers we choose to simplify the calculations

OpenStudy (anonymous):

Ok cool. Also, do you know how to find percentage error?

OpenStudy (anonymous):

you can say that ;)

OpenStudy (anonymous):

Haha how would you go about doing that?

OpenStudy (anonymous):

I would be so happy to do that

OpenStudy (anonymous):

I don't know how to find the percentage error

OpenStudy (anonymous):

ohh is that the last part on there?

OpenStudy (anonymous):

Haha my bad I got it all mixed up but I get it!

OpenStudy (anonymous):

okay, the percentage error is esimated with this rule:\[\frac{ f \prime (x) }{ f(x) }*100\]

OpenStudy (anonymous):

Cool, cool. Sorry got confused haha. Thanks for clearing that up

OpenStudy (anonymous):

okay but notice that I said esimated , and that what usually we need

OpenStudy (anonymous):

yup, it says to estimate for me so i think that will do it

OpenStudy (anonymous):

okay good luck and if you needed anything I will be around ;)

OpenStudy (anonymous):

You're a life saver right now

OpenStudy (anonymous):

Sorry I actually don't get why you had to do that last step. the .01-.0006

OpenStudy (anonymous):

Like I understand the math, but don't understand why you had to do it

OpenStudy (anonymous):

NEVERMIND IGNORE THAT! Brain fart

OpenStudy (anonymous):

I did nothing dude just substitution in the original equation which is \[y=y1+f \prime(x1) dx\]

OpenStudy (anonymous):

lol no no I liked that fart actually

OpenStudy (anonymous):

Lol the worst kind of fart

OpenStudy (anonymous):

okay there three things in mind right now: the first did you understand why we did the last step? second: is it safe to use that idiom ( brain fart) I kinda liked it lol the third: I have to admit, you are the coolest guy in this entire site ;)

OpenStudy (anonymous):

Hahah thanks bud! This was my first time using the site, so so far you have been my favorite! Yes I completely understand it now though, like I said my prof uses different terms so I just had to translate that, but then it was smooth. Lol hey whatever kinda farts you're into! I always get a kick out of brain fart :)

OpenStudy (anonymous):

Got to go take my test now that I'm feeling more confident about it. Thanks much for the help!

OpenStudy (anonymous):

lol , okay I wish you very good luck in your exam, just whenever you got stuck in a question, remember that question is nothing but a prof brain fart :)